OpenStudy (anonymous):

Model the function rule y = -2x - 3 with a table of values

OpenStudy (anonymous):

PLEASE DON'T LEAVE RUDE COMMENT OKAY THANK U

jimthompson5910 (jim_thompson5910):

Plug in x = 0 to get ... y = -2x-3 y = -2(0) - 3 y = 0 - 3 y = -3 So when x = 0, y is -3 Plug in x = 1 to get ... y = -2x-3 y = -2(1) - 3 y = -2 - 3 y = -5 So when x = 1, y is -5 Keep going with different various x values to build the table.

jimthompson5910 (jim_thompson5910):

If you plug in x = 2, what do you get?

OpenStudy (anonymous):

IAM SOOOOOOOOOOOO CLOSE OF GIVE UP RIGHT NOW

OpenStudy (unklerhaukus):

\[\times\]